OpenStudy (anonymous):

if you know that 7*8=56, how can you use the commutative property of multiplication to find the product of 8*7?

OpenStudy (sleepyhead314):

Hello and Welcome to OpenStudy! :) commutative property of multiplication is If a*b=b*a so if a*b=c then b*a=c